whats the square root of 10,000 times 2 times 50
Answer by checkley77(12569)

SQRT(10,000*2*50)=SQRT(1,000,000)=1,000 ANSWER.
OR SQRT(10,000)*2*50=100*100=10,000 ANSWER.